# Simulating Classroom Education with LLM-Empowered Agents

### v1

@@ Abstract
Large language models (LLMs) have been employed in various intelligent educational tasks to assist teaching. While preliminary explorations have focused on independent LLM-empowered agents for specific educational tasks, the potential for LLMs within a multi-agent collaborative framework to simulate a classroom with real user participation remains unexplored. In this work, we propose SimClass, a multi-agent classroom simulation framework involving user participation. We recognize representative class roles and introduce a novel class control mechanism for automatic classroom teaching, and conduct user experiments in two real-world courses. Utilizing the Flanders Interactive Analysis System and Community of Inquiry theoretical frame works from educational analysis, we demonstrate that LLMs can simulate traditional classroom interaction patterns effectively while enhancing user’s experience. We also observe emergent group behaviors among agents in SimClass, where agents collaborate to create enlivening interactions in classrooms to improve user learning process. We hope this work pioneers the application of LLM-empowered multi-agent systems in virtual classroom teaching.
@@
Zheyuan Zhang♠††thanks: Equal contribution, Daniel Zhang-Li♠11footnotemark: 1, Jifan Yu♡, Linlu Gong♠, Jinchang Zhou♠, Zhiyuan Liu♠, Lei Hou♠, Juanzi Li♠ ♠Department of Computer Science and Technology; ♡Institute of Education Tsinghua University, Beijing, 100084, China zheyuan-22@mails.tsinghua.edu.cn lijuanzi@tsinghua.edu.cn
@@ 1 Introduction
However, towards LLM-empowered multi-agent systems that involve real user participation, there are still several fundamental research questions that need to be explored. (1) Simulation Capability Assessment: To what extent can a multi-agent classroom powered by large models simulate real teacher-student interactions? (2) Learning Experience Measurement: Can students in such an intelligent teaching environment experience a high sense of presence and learn effectively? (3) Emergence Phenomenon Observation: What types of classroom behaviors may spontaneously arise in scenarios that integrate multiple agents?
In this work, responding to the questions above, we present SimClass, a Multi-Agent Classroom Simulation framework, and conduct real-world observation along with analysis based on it. To better simulate the classroom, we recognize representative class roles and design a novel class control mechanism with functional workflows. For systematic experiments, we deploy 2 different courses with prepared slides and teaching scripts as basis. 48 students are invited to join the classroom, learning and interacting with the system, and all the behavioral data is carefully recorded. Then we conduct experiments to explore the mentioned questions. (1) Firstly, we apply the Flanders Interaction Analysis System Amatari (2015) to evaluate the interactions happening in the SimClass and explore the interaction pattern of the agents’ classroom. (2) Secondly, we analyze the educational experience of these users, particularly with Community of Inquiry theory Garrison and Arbaugh (2007). (3) Finally, we summarize several emergent group actions during the experiment for qualitative analysis.
During our experiments, we observe the effectiveness of the class role and control mechanism design. Based on the problems we identified, experimental results show that: (1) Similarity: SimClass exhibits behaviors, interaction patterns, and characteristics similar to those of traditional classrooms; (2) Effectiveness: Multiple classroom agents enable users to engage more effectively in class and enhance their sense of presence; (3) Emergence: Our control mechanism spontaneously elicits the emergent behaviors in the multi-agent classroom system, including collaborative teaching and discussion, emotional company and discipline control. In summary, the LLM-based multi-agent system demonstrates the potential for simulating real classroom environments for educational purposes. We hope our work serves as a pioneering effort in this direction. The dataset of classroom interactions between users and multiple LLMs will be released soon for both education and AI researchers.
@@ 3.1 Overview
For the former concern, we categorize classroom interaction behaviors based on widely accepted pedagogy principles Schwanke (1981): Teaching and Initiation (TI), the teacher’s teaching and the feedback or ideas expressed by students; In-depth Discussion (ID), alignment, discussion, and multiple Q&A between teacher and students to help students construct understanding of concepts; Emotional Companionship (EC), encouraging students to learn, creating a positive learning atmosphere, and providing emotional support; and Classroom Management (CM), maintaining discipline, organizing disruptive behaviors, and guiding the classroom content. Given that these behaviors are realized through the varied Class Roles (denoted as , where each denotes a certain role), it is essential to ensure the diversity and coverage of proposed agents within the classroom.
For the latter concern, we need to ensure that the interactions among multiple agents within the system are finely and rhythmically controlled within the course content. Given the Learning Materials (denoted as , where each teaching script is organized by order), we propose a novel Session Controller to manage the course interaction flow based on class status and the help of a core manager agent Wu et al. (2023).
Based on these principles, we construct multiple class roles, implement class control, and ultimately derive the simulated classroom process.
@@ 3.2 Class Role Agentization
The teaching and learning process is presented as an informative, multi-round, and task-oriented communication Lave (1996). However, simply exchanging responses of LLMs inevitably faces significant challenges including role flipping, instruction repeating, and fake replies Qian et al. (2024). Consequently, following the classroom behaviors outlined previously, we define two types of agents: Teaching Agents and Classmate Agents. Each agent is facilitated through prompting LLMs and associated with one or more class roles, denoted as:
where is the role customization operation, is the system prompt with agent description.
@@
This type of agents are incorporated in addition to the teaching agents with distinct personality traits to better simulate traditional one-to-many classrooms, performing peer student roles. In this paper, we initialize typical classmates, while users can also freely customize and deploy more interesting classmate agents on the platform.
@@ 3.3 Classroom Session Controller
Unlike Standardized Operating Procedures (SOPs) multi-agent systems Qian et al. (2024); Hong et al. (2023), the classroom scenario is a dynamic group chat without a strict workflow, where agents need to dynamically determine the appropriate speaking timing. Therefore, we implement a controller that observes, makes decisions, and controls agents to behave based on the current Class State. The Session Controller includes the following modules: Class State Receptor, Function Executor, and Manager Agent.
Class State Receptor
Let the classroom dialogue history until time denote as , where is the utterance posted by agent or user (denoted as ). The class state is composed as:
@@
Functions
We design and divide the actions in the classroom into a functional hierarchy with two major categories. Tutoring functions can only be performed by teacher agent , such as teaching by displaying scripts and going to the next material page . Interacting functions can be performed by each agent . According to the context, the interaction will emerge as diverse classroom activities, which are discussed in subsequent experiments. These functions are pluggable, allowing the addition of newly defined functions for different agents, such as displaying exercises.
Manager Agent
Following AutoGen Wu et al. (2023) and MathVC Yue et al. (2024), we design a hidden and meta agent to regulate the speakers. This agent receives the current class state , observes and understands the class process, and decides the next action to be executed.The task of Manager Agent can be defined as:
@@ 3.4 Classroom Demonstration
After introducing the necessary component of the SimClass , we present the demonstration of an entire class process: (1) Initialization. At the beginning of the class, the first function will be executed, displaying the initial course script and slides. At this point, users can interact with the class, and the manager agent will start controlling the class flow; (2) Tutoring and Interaction: the manager agent will continuously observe and control the class based on the states, and other agents will perform diverse activities by collaboration. As the example shown in Figure 1, when a user asks about the course content, the classroom interaction flow may involve the assistant responding, the teacher providing additional information, and sometimes the classmate agents raising corresponding topics; (3) Ending. After all the learning materials are taught and the final discussion ends, the classroom will close and provide survey questions to users.
@@ 4 Experiments
To evaluate the performance of SimClass, we invite a group of university students to participate in the classroom to record interaction data and collect feedback from them. We also develop ablation systems to better understand the impact of various interaction types within SimClass. Our analyses mainly focus on three key aspects: classroom interactions, user experience, and the emergent group behaviors of the LLM-empowered agents.
@@ 4.1 Experimental Setup
Courses and Materials. We conduct experiments with two courses. Two experienced teachers are invited to design the slides and teaching scripts for the courses. The first one, TAGI, Towards Artificial General Intelligence, covers the development of AI and language models, consisting of 50 pages of slides, each with a corresponding teaching script. The second course, HSU, How to Study at University, addresses topics such as completing academic work, managing pressure, communicating with others, and achieving self-fulfillment, and includes 45 pages of slides and teaching scripts.
Systems. We use GPT-4 as the backbone LLM of both Class Roles and Manager Agent in SimClass. Besides, we implement two ablation systems to investigate the effects of different interaction types in the classroom. In the first system (SimClass w/o classmates), the classmate agents are removed, and only teacher agents are present. In the second system (SimClass w/o interactions), both classmate agents and user input are disabled, resulting in no interaction at all. The teacher can only conduct lectures persistently, and the Manager Agent is limited to the tutoring function.
Participants. In the experiment, we invite 48 university students from different majors to participate in learning two courses on SimClass, with each course and each setting involving 8 users. To ensure the quality control of course data, we invite course designers to create four questions for each course and require users to complete these questions after finishing the course. Data from participants with an accuracy of 50% or lower on these questions were excluded. Ultimately, data from 38 participants remained, with each course and setting having data from at least 5 users. Participants are informed that all course data is generated by AI and needed to be carefully discerned. Each participant receives the appropriate amount of compensation.
| Please rate the overall performance of the platform: |
| Cognitive Presence |
| Does the platform help students to understand concepts and master the corresponding knowledge? |
| Teaching Presence |
| Social Presence |
@@
Survey Content. In addition to the four test questions, each participant is required to complete a short survey composed of three questions regarding the experience after finishing the course. We apply the widely recognized Community of Inquiry (CoI) theory Garrison et al. (1999) in online learning to evaluate the experience of students. Specifically, we adapt the three key elements from CoI to measure the learning experience on SimClass: Cognitive Presence, the degree to which learners are able to construct and confirm meaning through sustained reflection and interaction; Teaching Presence, the extent to which the class is focused, designed, and planned with specific directions and learning objectives; and Social Presence, the ability of learners to project themselves socially and emotionally within a group Garrison and Arbaugh (2007). Students are asked to rate the system on a scale of [0,1,2], with a higher score indicating better performance according to detailed guidelines. The survey questions are listed in Table 1, and further details can be found in Appendix A.
4.2 Statistical Results
| Course | Teacher | Assistant | Students | User |
| TAGI | 353.0 | 82.3 | 123.0 | 18.9 |
| - w/o stu. | 358.2 | 71.1 | - | 13.9 |
| - w/o int. | 398.8 | - | - | - |
| HSU | 218.3 | 90.6 | 147.7 | 15.5 |
| - w/o stu. | 212.3 | 68.2 | - | 8.2 |
| - w/o int. | 228.5 | - | - | - |
Table 2 presents the average speech length of various roles and users across different settings. All systems employ the same teaching scripts, leading to the teacher’s speech being the longest and most closely aligned with the scripts. The assistant’s primary role is to maintain discipline, resulting in shorter dialogues. Classmate agents are generally more talkative, whereas users tend to use fewer words. Notably, the absence of classmate agents significantly reduces the speech length of users and assistant in both courses. The presence of classmate agents in the classroom appears to encourage users to engage in longer conversations.
4.3 Interaction Analysis
@@
To understand the dynamics of SimClass as a multi-agent classroom system, we encode classroom activities into quantitative behaviors. We utilize the Flanders Interaction Analysis System (FIAS) Amatari (2015), a valuable tool for analyzing the verbal behaviors in traditional classrooms. We adapt the method to our simulated classroom system, SimClass, where interactions occur in natural language.
@@
| Initiation | 9. Initiation | |
| Silence | Silence | 10. Silence or Confusion |
Encoding the Interactions. As shown in Table 3, the FIAS categorizes interactions into ten distinct types: seven for teachers, two for students, and one for silent. Labels represent Indirect Influence from the teacher, while labels indicate Direct Influence. When classroom activities are encoded as sequences, the proportion of each interaction type and their transitions can be decoded to reveal the classroom style, teaching style, and other features. For the classroom history of each student, we prompt GPT-4 to label interactions according to the ten communication categories. We assess the quality of GPT-4’s labeling in Appendix B. The classroom interactions are encoded as sequences, and the two-step transitions of classroom activities are recorded in a matrix . Following the method introduced by Amatari (2015), we add a (silence) to the beginning and end of each class sequence and sum the matrices of students in the same setting to provide a general view of the interactions: . To interpret the classroom interaction Matrix and observe features in SimClass, we report the following metrics designed by Amatari (2015):
@@
ID Ratio (IDR). This ratio measures the balance between a teacher’s indirect and direct methods of communication and teaching in the classroom. It is calculated by dividing the sum of tallies in categories (Indirect influence) by the sum of tallies in categories (Direct influence).
@@
| TAGI | 0.824 | 0.162 | 0.130 | 0.493 |
| - w/o stu. | 0.864 | 0.121 | 0.119 | 0.446 |
| - w/o int. | 0.980 | 0.000 | 0.005 | 0.000 |
| HSU | 0.847 | 0.136 | 0.127 | 0.597 |
| - w/o stu. | 0.896 | 0.087 | 0.164 | 0.233 |
| - w/o int. | 0.978 | 0.000 | 0.089 | 0.000 |
Results. Figure 2 presents the FIAS matrices of SimClass for TAGI and HSU courses. Each matrix is divided into four parts based on the type of interaction in the class, labeled as follows: A (top left): Interactions from teacher to teacher; B (top right): student or silence to teacher; C (bottom left): teacher to student or silence; and D (bottom right): student to student. The matrices reveal the following findings: (1) In part A, most teacher actions are associated with lecturing (Cat. ), where teachers primarily give lectures and interact with the class; (2) Part B demonstrates the teacher’s responses to students. When students initiate ideas or responses to teachers, the teacher praises (Cat.), accepts their ideas (Cat.), or continues teaching; (3) Part C shows student actions in response to teachers, where students mostly initiate questions or respond to lectures; (4) Part D shows that student-to-student interactions and discussions occur periodically. The results of the ablation systems, demonstrated in Appendix C, indicate that interactions are much less diverse after ablation.
Table 4 presents the metric results of FIAS. TT and ST represent the proportion of teacher and student speaking time, respectively. In traditional classrooms, the statistics for TT and ST range from 77.1%-82.6% and 17.4%-22.9% (excluding silence) Zhang et al. (2023). SimClass exhibits a similar distribution. The IDR is low, which is partly due to the higher proportion of script-based teaching. The SIR is relatively high, especially in scenarios involving classmate agents, where there are more instances of students initiating questions.
Conclusion of Interactive Analysis. From the perspective of ratio analysis (Table 4) and interaction distribution (Figure 2), SimClass demonstrates the characteristics of the traditional classroom, effectively simulates traditional classrooms and has the potential to achieve the performance of real classrooms. We further investigate the user experience and illustrate a few in the following sections.
4.4 User Experience

### v2

@@ Abstract
Large language models (LLMs) have been applied across various intelligent educational tasks to assist teaching. While preliminary studies have focused on task-specific, independent LLM-empowered agents, the potential of LLMs within a multi-agent collaborative framework for classroom simulation with real user participation remains unexplored. In this work, we propose SimClass, a multi-agent classroom simulation teaching framework. We recognize representative class roles and introduce a novel class control mechanism for automatic classroom teaching, and conduct user experiments in two real-world courses. Using the Flanders Interactive Analysis System and Community of Inquiry theoretical frameworks from educational analysis, we demonstrate that LLMs can simulate a dynamic learning environment for users with active teacher-student and student-student interactions. We also observe group behaviors among agents in SimClass, where agents collaborate to create enlivening interactions in classrooms to improve user learning process. We hope this work pioneers the application of LLM-empowered multi-agent systems in virtual classroom teaching.
@@
Zheyuan Zhang♠††thanks: Equal contribution, Daniel Zhang-Li♠11footnotemark: 1, Jifan Yu♡, Linlu Gong♠, Jinchang Zhou♠, Zhanxin Hao♡, Jianxiao Jiang♡, Jie Cao♡, Huiqin Liu♡, Zhiyuan Liu♠, Lei Hou♠, Juanzi Li♠ ♠Department of Computer Science and Technology; ♡Institute of Education Tsinghua University, Beijing, 100084, China zheyuan-22@mails.tsinghua.edu.cn lijuanzi@tsinghua.edu.cn
@@ 1 Introduction
However, several fundamental research questions for LLM-empowered multi-agent systems with real user participation remain: (1) Simulation Performance: How well can a multi-agent classroom simulate real-time teacher-student interactions? (2) Learning Experience: Can students in such environment experience a high sense of presence and learn effectively? (3) Group Behavior Observation: What behaviors may arise spontaneously in multi-agent scenarios?
In response to the research questions above, we present SimClass, a multi-agent classroom simulation framework, and conduct real-world experiments and analysis on it. For better simulation, we identify representative class roles and design a novel class control mechanism (Figure 1). We deploy two different courses with prepared slides and teaching scripts as the foundation. We conduct online experiments with over 400 students, who participated in the courses and interacted with the system, with all the behavioral data carefully recorded. Additionally, we constructed ablation systems and invited another 48 students for further experiments. Our research addresses the following parts: (1) We apply the Flanders Interaction Analysis System Amatari (2015) to evaluate the interactions within the SimClass and examine the teaching style of the agents. (2) We analyze the learning outcomes and educational experiences of these users, using Community of Inquiry theory Garrison and Arbaugh (2007). (3) Lastly, we identify group behaviors of agents for qualitative analysis.
The experimental results demonstrate the effectiveness of the class roles and control mechanism design in the following aspects: (1) Performance: SimClass fosters a vivid learning environments with lively teacher-student and student-student interactions; (2) Experience: Students retain the knowledge gained in the SimClass, with increased interactions contributing to improved learning outcomes. The presence of multiple classroom agents enhances user engagement and strengthens their sense of presence; (3) Behavior: The control mechanism elicits group behaviors in the multi-agent classroom, including collaborative teaching, discussions, emotional company and discipline management. In summary, the LLM-based multi-agent system shows great potential for simulating classroom for educational purposes. We hope our work serves as a pioneering effort in this area. The dataset of classroom interactions will be released soon for both education and AI researchers.
@@ 3.1 Overview
For the former concern, we categorize classroom interaction behaviors based on widely accepted pedagogy principles Schwanke (1981): Teaching and Initiation (TI), the teacher’s teaching and the students’ feedback or ideas; In-depth Discussion (ID), alignment, discussion, and multiple Q&A to help students construct understanding of concepts; Emotional Companionship (EC), encouraging students to learn, creating a positive learning atmosphere, and providing emotional support; and Classroom Management (CM), maintaining discipline, organizing disruptive behaviors, and guiding the classroom content. Given that these behaviors are realized through the varied Class Roles (denoted as , where each denotes a certain role), it is essential to ensure the diversity and coverage of proposed agents within the classroom.
For the latter concern, we need to ensure that the interactions among multiple agents within the system are finely and rhythmically controlled within the course content. As shown in Figure 1, given the Class Roles and Learning Materials (denoted as , where each teaching script is organized by order), we propose a novel Session Controller to manage the course interaction flow based on class status and the help of a core manager agent Wu et al. (2023).
Based on these principles, we construct multiple class roles, implement class control, and ultimately derive the simulated classroom process. Their prompts are shown in Appendix.
@@ 3.2 Class Role Agentization
The teaching process is presented as an informative, multi-round, and task-oriented communication Lave (1996). However, simply exchanging responses of LLMs inevitably faces significant challenges including role flipping, instruction repeating, and fake replies Qian et al. (2024). Consequently, following the classroom behaviors outlined previously, we define two types of agents: Teaching Agents and Classmate Agents. Each agent is facilitated through prompting LLMs and associated with one or more class roles, denoted as:
where is the role customization operation, is the system prompt with agent description. All roles designs and corresponding prompts were crafted with input from experienced teaching practitioners. Relevant technologies, such as question generation Kurdi et al. (2020) and retrieval-augmented generation Lewis et al. (2020), can also be integrated into the construction of class roles.
@@
This type of agents are incorporated in addition to the teaching agents with distinct personality traits to perform peer student roles. In this paper, we initialize typical classmates, while users can also freely customize more interesting classmate agents on the platform.
@@ 3.3 Classroom Session Controller
Unlike multi-agent systems with Standardized Operating Procedures (SOPs) Qian et al. (2024); Hong et al. (2023), the classroom scenario is a dynamic group chat without a strict workflow, requiring agents to determine appropriate speaking times on the fly. Therefore, we implement a Session Controller that observes, decides, and directs agent behavior based on the current Class State. It comprises three modules: Class State Receptor, Function Executor, and Manager Agent.
Class State Receptor. Let the classroom dialogue history until time denote as , where is the utterance posted by agent or user (denoted as ). The class state is composed as:
@@
Functions. We design and divide the actions in the classroom into a functional hierarchy with two major categories. Tutoring functions can only be performed by teacher agent , such as teaching by displaying scripts and going to the next material page . Interacting functions can be performed by each agent . According to the context, the interaction will emerge as diverse classroom activities, which are discussed in subsequent experiments. These functions are pluggable, allowing the addition of newly defined functions for different agents, such as displaying exercises.
Manager Agent. Following AutoGen Wu et al. (2023) and MathVC Yue et al. (2024), we design a hidden and meta agent to regulate the speakers. This agent receives the current class state , observes and understands the class process, and decides the next action to be executed. The task of Manager Agent can be defined as:
@@ 3.4 Classroom Demonstration
After introducing the necessary components of the SimClass, we demonstrate a complete class process: (1) Initialization. At the start, the first function executes, displaying the initial script and slides. Users can begin interacting, and the manager agent takes control of the class flow; (2) Tutoring and Interaction: the manager agent will continuously observe and control the class based on the states, selecting appropriate functions and speakers, and coordinates agent collaboration. As shown in Figure 1, when a user asks about the content, the classroom interaction flow may involve the assistant responding, the teacher adding details, and sometimes the classmate agents raising relevant topics; (3) Ending. After all the learning materials are taught and the final discussion ends, the classroom will close and provide quizzes to users.
@@ 4 Experiments
We focus on three research questions to evaluate and understand SimClass as a multi-agent learning environment: (1) What is the performance of SimClass? (2) What are the impacts of various interaction types within SimClass (e.g. the roles of student agents)? (3) How do agents behave in SimClass? To address these questions, we deploy SimClass online and invite a group of university students to use the system. With the Institutional Review Board (IRB) approval from our institution, we design quizzes and collected interaction data for further analysis. Additionally, we develop ablation systems to investigate the influence of class roles and interactions in the environment.
@@ 4.1 Experimental Setup
Courses and Materials. We conduct experiments with two courses. The first, TAGI, Towards Artificial General Intelligence, covers AI development and language models across six meticulously designed chapters. The second, HSU, How to Study at University, focuses on academic skills, stress management, comminication, and self-fulfillment, spanning seven well-structured chapters. While both courses contain structured slides and teaching scripts, TAGI emphasizes knowledge acquisition, and HSU aims at skill development.
Systems. We use GLM-4 GLM et al. (2024) as the backbone LLM for both Class Roles and Manager Agent due to cost and concurrency constraints in the online system. To explore the impact of class roles, we deploy three ablation systems with the model replaced by GPT-4V. The first replicates the original system, the second removes classmate agents (retaining only the teacher and assistant), and the third disables both classmate agents and user input, with the teacher conducting uninterrupted lectures LLMs can effectively deliver courses without modifying agent prompts.
Participants. We recruited over 400 university students from various majors for the online learning system, with 118 completing all of the chapters (77 in TAGI, 41 in HSU). An additional 48 students participated in the ablation study for only the first course chapter. To ensure data quality of ablation systems, students took a brief test after the course, assessing whether they used the system and remember some of the basic concepts covered, and data from those scoring below 50% was excluded. Participants were informed that course content is AI-generated, and their interaction data would be only used for research. All participants were compensated above the national average hourly wage.
Data Collection. The data collected for the online and the ablation systems emphasizes different aspects to address specific research questions:
(1) Online System: We meticulously recorded all user interactions for interaction analysis (Section 4.2.1). To evaluate students’ learning outcomes, we invited educational practitioners to design quizzes after each chapter and a final exam for the knowledge acquisition course, TAGI, to assess knowledge retention (Section 4.2.2). For the practical course, HSU, we employed a self-reported method, where students wrote self-summaries.
@@
(2) Ablation systems: We examined how class roles affect learning by tracking interactions and developing a brief survey based on the widely recognized Community of Inquiry (CoI) theory Garrison et al. (1999). The survey measures three elements: Cognitive Presence, the degree to which learners are able to construct and confirm meaning through sustained reflection and interaction; Teaching Presence, the extent to which the class is focused, designed, and planned with specific learning objectives; and Social Presence, the ability of learners to project themselves socially and emotionally within a group Garrison and Arbaugh (2007). Following prior research Yu et al. (2022); Tu et al. (2023), students rated the system on a [0,1,2] scale according to detailed guidelines, with higher scores indicating better performance. Survey questions and guidelines are detailed in Appendix B.
4.2 Online System Results
We demonstrate the performance of SimClass in the online system by analyzing student interactions and their learning outcomes.
4.2.1 Interaction Analysis
@@
To understand the dynamics of SimClass as a multi-agent classroom system, we encode classroom activities into quantitative behaviors. We employ the Flanders Interaction Analysis System (FIAS) Amatari (2015), a widely adopted tool for analyzing verbal behaviors in traditional classrooms. We adapt the method to our simulated classroom system, where interactions occur in natural language, to investigate the teaching patterns of SimClass.
@@
| Initiation | 9. Initiation |
Encoding the Interactions. As shown in Table 1, the FIAS categorizes interactions into nine distinct types: seven for teachers and two for students (we exclude the Silence category from the original FIAS, as it is not applicable to SimClass due to the real-time responses of models and the difficulty of defining or detecting silence in online systems). Labels represent Indirect Influence from the teacher, while labels indicate Direct Influence. For the classroom history of each student, we prompt GPT-4 to label interactions according to the nine categories, and assess the quality of labeling in Appendix A. The classroom interactions are encoded as sequences, and the two-step transitions of classroom activities are recorded in a matrix . To interpret the Matrix and observe features in SimClass, we report the following frequently used metrics:
@@
ID Ratio (IDR). ID Ratio measures the balance between a teacher’s indirect and direct methods of teaching in the classroom. It is calculated by dividing the sum of tallies in categories (Indirect) by the sum of tallies in categories (Direct).
@@
Results. We randomly sampled 10 students each who completed the courses and summed their matrices to view the interactions: . Figure 2 presents the FIAS matrices of SimClass for TAGI and HSU courses, with each divided into four parts based on the type of classroom interaction, labeled as follows: A (top left), Teacher Lecturing: Most teacher actions involve lecturing (Cat. ), where the teacher primarily delivers lessons and interacts with the class. B (top right), Student to Teacher: It demonstrates the teacher’s responses to students. When students initiate ideas or responses to teachers, the teacher praises (Cat.), accepts their ideas (Cat.), or continues teaching. C (bottom left), Teacher to Student: It highlights student responses to the teacher, where students frequently ask questions or react to lectures, reflecting the active participation in the classroom. D (bottom right), Student to Student Interactions: This part shows that student-to-student discussions occur periodically. Overall, the classroom exhibits frequent interactions, both between the teacher and students, as well as among the students themselves..
@@
| TAGI | 0.816 | 0.184 | 0.058 | 0.896 |
| HSU | 0.863 | 0.137 | 0.124 | 0.917 |
| ET | 0.771 | 0.229 | 1.473 | 0.121 |
| NT | 0.826 | 0.174 | 0.885 | 0.106 |
Table 2 presents the metric results of FIAS, illustrating the teaching style of SimClass. When compared with human classrooms reported by Zhang et al. (2023), the TT and ST ratios are similar, indicating a comparable speaking balance between SimClass and traditional classrooms (with silence removed from both scenarios for a fair comparison). The IDR is relatively low, partly due to the higher proportion of script-based teaching. Meanwhile, the SIR, which reflects the proportion of student-initiated interactions, is relatively high, suggesting a more democratic online learning environment where students feel comfortable asking questions and expressing themselves.
In summary, SimClass fosters a dynamic learning environment with active interactions between teachers and students, as well as among students themselves. Compared to traditional classrooms, students in SimClass are more proactive in initiating discussions and expressing their ideas.
4.2.2 Learning Outcome Analysis
We assess students’ learning outcomes in the online system through quizzes or self reports. In TAGI, all the questions in the quizzes are multiple-choice, with one or more correct answers, requiring all correct answers for full marks. The final exam draws from previous quizzes to evaluate knowledge retention. The average quiz scores (with full mark as 1) is presented in Table 3. The final exam score aligns with the average quiz scores (0.68), indicating students’ consistent retention of the material.
| Quiz | Final | |||||
| 1st | 2nd | 3rd | 4th | 5th | 6th | - |
| 0.64 | 0.53 | 0.66 | 0.82 | 0.78 | 0.66 | 0.65 |
We further analyzed the quiz results across different students. As shown in Figure 3, the scatter plot reveals a clear positive correlation between normalized quiz scores and both the log-transformed message length and message number, with a Pearson correlation coefficient of and , which is both statistically significant (). These findings suggest that students who engage more actively — by sending more and longer messages — tend to achieve higher average quiz scores. We further demonstrate the results of self-report surveys for HSU in Appendix C.2.
4.3 Ablation System Results
We investigate the impacts of various interaction types within SimClass through statistical results and the CoI outcomes from the ablation systems.
4.3.1 Statistical Results
Table 4 presents the average speech length of teacher and users across different settings in the ablation systems. As all systems employ the same teaching scripts, the teacher’s speeches are largely similar, with slight variations during instant interactions. Notably, removing classmate agents significantly reduces users’ speech length in both courses, whose presence encourages longer user conversations. Additional results of other agents and ablation systems are provided in Appendix C.3.
| Course | Teacher | Assistant | User |
| TAGI | 353.0 | 82.3 | 18.9 |
| - w/o cla. | 358.2 | 71.1 | 13.9 ( 26.5%) |
| HSU | 218.3 | 90.6 | 15.5 |
| - w/o cla. | 212.3 | 68.2 | 8.2 ( 45.2%) |
4.3.2 Community of Inquiry Analysis